OpenStudy (anonymous):

sec^2(50) - 4 + tan^2(50).. is the answer to this - 3

Directrix (directrix):

1 + tan^2(50) = sec^2(50) ---> trig identity ------------------- sec^2(50) - 4 + tan^2(50) = 1 + tan^2(50) - 4 + tan^2(50) = 1 - 4 = -3 Yes, you are correct.

OpenStudy (anonymous):

Thank you i got the same answer but i was just trying to see if i was right

OpenStudy (anonymous):

okay so what about this problem given sec (theta) = 3 what is the exact value of csc (90 - (theta))

Directrix (directrix):

Secant and Cosecant are co-functions which means that they have the same value at complementary angles. A case in point for sine and cosine would be that: sin(30) = cos(60) So, csc (90 - (theta)) = 3. Agree?

OpenStudy (anonymous):

yeah someone just explained it abit for me i was still confused for a bit but accepted the statement to be true the problem comes when its cos (theta) = 1/6 and you have to find sin(90 - theta) and csc( pie/2 - theta)

Directrix (directrix):

Would you write the question in a new post along with the instructions? (You need these trig identities to do this stuff. :) )

Directrix (directrix):

Thanks for doing that.